OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer? Numerical Answers Expected!

OpenStudy (anonymous):

thanx i got more to igbasallote , ill keep giving adwards

OpenStudy (lgbasallote):

x = first integer x+ 1 = 2nd x+2 = 3rd x + x + 1 + x + 2 = 267 3x + 3 = 267 3x = 264 x = 88 so largest is x+2 = 90 btw...it's Lgbasallote